7 Electric forces and pairs of charges Two charged particles, A and B, are held fixed. Let FA on B be the magnitude of the force exerted by A on B and FB on A be the magnitude of the force exerted by B on A. Which of the following is true? Explain your answer +3.0C +12.0C i) FB on A FA on B i) FB on A FA on B ii) FB on A 3FA onB iv) FB on A 4FA onB v) FB on A 12FA on B
Answer #1

Force due to the charged particle F= K*Q1*Q2/r^2

FB on A = K*12*3*/r^2 = K*36/r^

FA on B = K*3*12/r^2 = K*36/r^2

So

FB on A =FA on B

Option ii is true